### v1

# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

Abstract
Large Multimodal Models (LMMs) have shown significant reasoning capabilities by connecting a visual encoder and a large language model. LMMs typically use a fixed amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content. Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which increase the number of visual tokens significantly. However, due to the design of the Transformer architecture, computational costs associated with these models tend to increase quadratically with the number of input tokens. To tackle this problem, we explore a token reduction mechanism and find, similar to prior work, that many visual tokens are spatially redundant. Based on this, we propose PruMerge, a novel adaptive visual token reduction approach, which largely reduces the number of visual tokens while maintaining comparable model performance. We first select the unpruned visual tokens based on their similarity to class tokens and spatial tokens. We then cluster the pruned tokens based on key similarity and merge the clustered tokens with the unpruned tokens to supplement their information. Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14.4 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks. Code and checkpoints will be released.
1 Introduction
Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities. LLMs are usually high-capacity Transformers pretrained with a large-scale text corpus. Large Multimodal Models (LMMs), inherit the LLMs for text generation, but also leverage a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
LMMs need substantial computation to conduct inference. The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM. For example, the commonly used CLIP visual encoder, ViT-L, only has 0.3B parameters, while the corresponding LLM such as LLaMA [Touvron et al., 2023] or Vicuna [Vicuna, 2023] can have 7B or 13B parameters. As a result, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023]. However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap in visual question-answering and reasoning such as VQAv2 and MM-Bench [Chu et al., 2024]. A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation. The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens. For example, in LLaVA-1.5, 576 visual tokens are appended, leading to high training and inference costs. Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
Refer to caption (a) Main idea of LLaVA-PruMerge. Refer to caption (b) PruMerged Token Visualization.
Figure 1: (a) We prune and merge the visual tokens coming from the vision encoder, while keeping all other procedures of the LMM the same. By reducing the number of visual tokens, our proposed method, PruMerge, significantly reduces the computation cost for text generation in LMMs. (b) The visualizations of the attentive tokens. We design a token reduction method to adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently. More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images. Besides, such attentive tokens are usually located at the regions with dense information.
In our study, we find that such visual tokens are redundant, similar to findings in previous related work Bolya et al. [2023], Liu et al. [2022], and most of the visual tokens can be pruned without largely sacrificing the performance. In particular, we find that the activations are very sparse upon the similarity matrix between the class token and spatial patches, which indicates that only a small amount of the visual tokens are related to the key visual information in the image. Motivated by this, we use this similarity to select important visual tokens. Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune the visual tokens. Moreover, we merge the visual tokens using $k$-nearest neighbor and update the sampled visual tokens via weighted averaging, which further enhances performance. Finally, we optionally finetune the LLM to let the model better adapt to our token deduction design.
Empirically, LLaVA-PruMerge can effectively and adaptively prune the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 6.9% of visual tokens, which is around 40 tokens on average, our model can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Our work demonstrates the effectiveness of building efficient large multimodal models from the perspective of visual token pruning and paves the road for further research.
2 Related Work
2.1 Large Multimodal Models (LMMs)
Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text. Large Multimodal Models (LMMs) [Liu et al., 2023b, Zhu et al., 2023, Yin et al., 2023, Zhang et al., 2024] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format. More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023]. Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022]. The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a]. In this paper, we reduce the number of visual tokens by leveraging the similarity between the class token and the spatial patch tokens.
2.2 Efficient LMMs
The need for cross-modal capabilities in resource-limited scenarios has become increasingly important. Despite advancements in LMMs, their large-scale training and deployment incur significant computational costs, necessitating efficient parallel device implementations. Google’s Gemini [Team et al., 2023] is a leader in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices. However, Gemini remains closed-source. Open-source initiatives, like LLaVA-1.5 [Liu et al., 2023a], utilize advanced compression techniques, such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024]. Further efforts towards efficient LMMs include MobileVLM [Chu et al., 2023], which develops a compact LLM and an efficient multimodal feature projector, and its successor, MobileVLM-v2 [Chu et al., 2024], which explores improved training strategies for mobile scenarios. TinyGPT-V [Yuan et al., 2023a] leverages the advanced Phi-2 [Javaheripi et al., 2023] LLM to surpass the performance of significantly larger models. Similarly, LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for broader generalizability. TinyLLaVA [Zhou et al., 2024] investigates the impacts of architectural choices, data quality, and training strategies, demonstrating that smaller LMMs can match the performance of their larger counterparts with optimized data and training. MoE-LLaVA [Lin et al., 2024] adapts Mixture of Experts (MoE) to mitigate model degradation due to sparsity.
2.3 Token Reduction
The notorious squared complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the sequence length. Sparse attention such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context. Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching. In recent LMMs [Liu et al., 2023b, Zhu et al., 2023], prefix visual tokens serve as a fixed budget for context, which becomes one of the leading factors for their efficiency. In our study, we find that by simply pruning and merging visual tokens based on their similarity, we can achieve comparable performance using less than one tenth of the original tokens.
3 Method: Token Pru-Merging
In this section, we first review the basic implementation of large mutilmodal models (LMMs), with a particular focus on the visual encoder component (i.e., Vision Transformer). We highlight the direct correlation between the number of visual tokens and the efficiency of LMMs (Sec. 3.1). Next, we present a plug-and-play token reduction method specifically designed for LMMs, called token PruMerge. Our method features two key components: (1) Adaptive Important Token Selection (AITS) via Outlier Detection which adaptively determines the optimal number of visual tokens to retain based on the unique characteristics of the image (Sec. 3.2); and (2) Token Supplement (TS) via Similar Key Clustering, which facilitates more efficient processing without compromising the model’s performance by maintaining the integrity and richness of the visual information (Sec. 3.3).
3.1 Preliminaries
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. ViTs perform tokenization by dividing an input image into patches and projecting each patch into a token embedding. An extra class token [CLS] is added to the set of image tokens and is responsible for aggregating global image information and performing classification. Let $n$ denote the number of input tokens (image patches) to the ViT. ViTs are composed of a series of transformer blocks. An assembly of several key components composes each transformer block: a multi-head self-attention layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016], all of which work together to enhance the model’s ability to capture and interpret visual information [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: the query vector $\mathbf{q}$, the key vector $\mathbf{k}$, and the value vector $\mathbf{v}$, utilizing three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. The self-attention mechanism computes the relevance of each item to other items as follows:
$\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (3.1)
where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ represents the dimension of $\mathbf{q},\mathbf{k},\mathbf{v}$. In the last encoder layer of ViT, the [CLS] token is used for classification. Similarly, the attention between [CLS] token and other visual tokens are performed via the attention mechanism:
$\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (3.2)
The multi-head self-attention framework allows for simultaneous focus on multiple positions, offering diverse representation subspaces. This is achieved by employing distinct query, key, and value matrices for different heads, which project the input vectors into varied representation subspaces. Subsequent to the self-attention layers is the feed-forward network (FFN). This network consists of two linear transformation layers separated by a nonlinear activation function and is represented as:
$\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3.3)
with $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ being the parameter matrices of the linear transformation layers and $\sigma$ denoting the nonlinear activation function, such as GELU [Hendrycks & Gimpel, 2016]. The general forward pass of ViT is illustrated in the left part of Figure 2.
Figure 2: The conceptual idea of LLaVA-PruMerge. Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens; (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster. Here $m$ denotes the visual token compression ratio.
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a series of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which is responsible for mapping the encoded features from other modalities $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of LMMs is depicted in Figure 1. It is important to note that the computational cost associated with these models tends to increase quadratically with the number of input tokens fed into the LLM backbone [Tay et al., 2022]. There is an increasing demand for the processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
3.2 Adaptive Important Token Selection via Outlier Detection
Figure 3: The distribution of attention values between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token.
The most straight-forward solution to improve the efficiency of visual token utilization in LMMs is to prune the redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022]. To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
As discussed in Sec. 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information. On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [Radford et al., 2021] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token. To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2. Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3. This sparse distribution underpins our methodology for identifying crucial visual tokens. By employing outlier detection algorithms, we aim to adaptively select visual tokens that best represent an image’s features effectively.
Interquartile Range (IQR) Method for outlier detection. To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [Boukerche et al., 2020], a statistical technique known for its robustness in outlier detection. The essence of the IQR method lies in its capability to establish a boundary or “fence” that delineates the normal range of data. This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1. Specifically, the computation is as follows: the “lower fence” is set at 1.5 $\times$ IQR below Q1, and the “upper fence” is set at 1.5 $\times$ IQR above Q3. Any attention values residing outside these fences are classified as outliers. Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context.
As shown in Figure 0(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image. Simpler images such as ”Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens. (2) The sampled tokens are spatially aligned with the important details, while less tokens are assigned for the background. Such visualizations align with our visual token sampling design.
Algorithm 1 Token PruMerge algorithm for reducing the number of visual tokens.
1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$. The penultimate layer’s output tokens, $\mathbf{Y}=\{\mathbf{y}_{1},\cdots\mathbf{y}_{n}\}$. $n$ is the number of input visual tokens.
2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
3:Token PruMerge:
4:Calculate attention between visual token and class token $\mathbf{a}_{[cls]}$ using Equation 3.2.
5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ index $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
6:for $p=\{i_{1},\cdots,i_{m}\}$ do
7:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
8:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
9:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
10:end for
11:(see Sec. 3.3).
12:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
3.3 Token Supplement via Similar Key Clustering
Following the selection of more informative visual tokens, we next optimize the utilization of the remaining tokens. While pruned tokens may initially seem extraneous, they hold potential value for the perception capabilities of the LLM backbone. This potential arises particularly in cases where an image contains large object parts that dominate the scene. In such scenarios, overly aggressive pruning could inadvertently diminish the model’s ability to represent the image comprehensively.
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Luckily, transformers natively solve this problem with QKV self-attention. Specifically, the keys (K) already summarize the information contained in each token for use in dot product similarity. Thus, we use a dot product similarity metric (e.g., cosine similarity) between the keys of each token to determine which tokens contain similar visual information [Bolya et al., 2023]. Specifically, the similarity between visual tokens is computed as
$\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (3.4)
which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorization form for the set of all tokens ${1,2,\cdots,n}$.
With the similarities between visual tokens established, we simply find the $k$-nearest neighbors for each unpruned token, which act as the cluster centers. The integration of pruned tokens into these clusters is guided by their respective class attentions $\mathbf{a}[i]$, enabling a refined representation of each unpruned token through a weighted sum. This procedure is outlined in Algorithm 1.
4 Experiments
We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then demonstrate the effectiveness of each component in our model in Sec 4.3.
Table 1: Comparison with large multimodal models on six benchmarks. Our proposed approach LLaVA-PruMerge can adaptively prune and merge visual tokens, which uses only 6.9% visual tokens on average and achieves competitive performance compared to the original LLaVA-1.5 model.
Method LLM Res. PT IT VQAv2 SQAI VQAT POPE MME MMB
BLIP-2 Vicuna-13B 224 129M - 41.0 61 42.5 85.3 1293.8 -
InstructBLIP Vicuna-7B 224 129M 1.2M - 60.5 50.1 - - 36
InstructBLIP Vicuna-13B 224 129M 1.2M - 63.1 50.7 78.9 1212.8 -
Shikra Vicuna-13B 224 600K 5.5M 77.4 - - - - 58.8
IDEFICS-9B LLaMA-7B 224 353M 1M 50.9 - 25.9 - - 48.2
IDEFICS-80B LLaMA-65B 224 353M 1M 60.0 - 30.9 - - 54.5
Qwen-VL Qwen-7B 448 1.4B 50M 78.8 67.1 63.8 - - 38.2
Qwen-VL-Chat Qwen-7B 448 1.4B 50M 78.2 68.2 61.5 - 1487.5 60.6
LLaVA-1.5 Vicuna-7B 336 558K 665K 78.5 66.8 58.2 85.9 1510.7 64.3
LLaVA-1.5 + Ours Vicuna-7B 336 558K 665K 72.0 68.5 56.0 86.3 1350.3 60.9
LLaVA-1.5 Vicuna-13B 336 558K 665K 80.0 71.6 61.3 85.9 1531.3 67.7
LLaVA-1.5 + Ours Vicuna-13B 336 558K 665K 72.8 71.0 58.4 86.2 1428.2 62.3
4.1 Main Results
We apply our method to LLaVA-1.5 [Liu et al., 2023a], a recent state-of-the-art LMM. We further finetune LLaVA-1.5 using LoRA [Hu et al., 2022] for 1 epoch using the LLaVA-1.5 instruction fine-tuning data [Liu et al., 2023a] with our reduced visual tokens.
We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c]. As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 while performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023]. Specifically, in POPE and SQA, our approach even shows better performance than LLaVA-1.5.
4.2 Efficiency Analysis
To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis as developed in [Yuan et al., 2024]. Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs. Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens. Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis. Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times to match the token count of the text prompt ($576/14.4\approx 40$). The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process. Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
Table 2: Computation Cost Analysis. The development device is TESLA V100 GPU, and time estimated by the roofline model represents the theoretical performance that the hardware can achieve.
Method LLM Quantization FLOPs Prefill Total Storing
Backbone (T) Time (ms) Memory (G) Activation (G)
LLaVA-1.5 Vicuna-7B FP16 9.3 88.6 23.3 4.60
LLaVA-1.5 w/ PruMerge Vicuna-7B FP16 0.91 15.3 13.7 0.28
LLaVA-1.5 Vicuna-7B INT4 2.3 151.6 5.9 1.20
LLaVA-1.5 w/ PruMerge Vicuna-7B INT4 0.28 14.9 3.5 0.07
LLaVA-1.5 Vicuna-13B FP16 18.2 170.5 41.6 7.30
LLaVA-1.5 w/ PruMerge Vicuna-13B FP16 1.80 29.5 26.6 0.44
LLaVA-1.5 Vicuna-13B INT4 4.6 294.9 10.5 1.80
LLaVA-1.5 w/ PruMerge Vicuna-13B INT4 0.45 29.0 6.8 0.11
It is important to emphasize that the benefits of PruMerge  extend beyond mere efficiency gains. Our token reduction strategy can complement other LLM acceleration techniques, such as quantization and factorization [Yuan et al., 2023b]. This orthogonal relationship underscores the versatile potential of PruMerge to contribute to a broader spectrum of efficiency-enhancing strategies.
4.3 Ablation Study
4.3.1 Token Sampling Strategy Analysis
Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec. 3.2), which results in 40 visual tokens per image on average. The visualization is shown in Figure 4 (b).
Sequential sampling: We sample $N$ tokens in the flatted visual tokens, shown in Figure 4 (c). Here the first 40 tokens are sampled for an apple-to-apple comparison.
Spatial sampling: Sampled $N$ tokens are spatially distributed evenly across the image, as shown in Figure 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (36 tokens), and 7 $\times$ 6 (42 tokens).
Figure 4: The visualization of the different visual tokens sampling strategy.
Note that all the experiments are done via the training-free manner. As shown in Table 3, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge. Besides, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019]. Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details. This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 0(a) (b), where more attentive tokens are distributed on the foreground text in the images.
Table 3: Ablation on training free and fine-tuning for our approach. With further fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
Approach Number of Visual Tokens SQA TextVQA POPE MME
Sequential 40 64.60 42.72 13.2 703.60
Spatial $6\times 6=36$ 66.58 46.84 67.9 1169.10
$5\times 8=40$ 66.19 46.85 70.4 1180.23
$8\times 5=40$ 68.12 47.42 71.1 1142.32
$6\times 7=42$ 67.08 47.96 72.5 1220.89
$7\times 6=42$ 66.53 47.49 71.6 1199.10
LLaVA-PruMerge 40 68.07 54.00 76.2 1250.07
4.3.2 Effectiveness of Each Module in PruMerge
Table 4: Ablation Studies for Adaptive Important Token Selection (AITS, Sec. 3.2) and Token Supplement (TS, Sec. 3.3). With these modules, the downstream performance can be progressively improved.
Method LLM SQAI VQAT POPE MME
LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 85.7 1221.6
LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 86.3 1350.3
Here, we study the effectiveness of each module in our design based on LLaVA-1.5. Note that we maintain the same amount of visual tokens (6.9%, 40 tokens) across all settings. As shown in Table 4, after progressively adding the proposed modules, including Adaptive Important Token Selection (AITS) and Token Supplement (TS), the downstream performance can be further enhanced.
4.3.3 Training Analysis: Training-free v.s. Fine-tuning
LLaVA-PruMerge can be conducted both in training-free and training-needed manners. With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks. As shown in Table 5, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].
Table 5: Ablation on training free and fine-tuning for our approach. With further fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
Method LLM SQAI VQAT POPE MME
LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
LLaVA-PruMerge w.o. fine-tuning Vicuna-7B 68.0 54.0 76.2 1250.1
LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 86.3 1350.3
5 Conclusion
In this paper, we improve the efficiency of Large Multimodal Models (LMMs) from the perspective of reducing the quantity of visual tokens. By leveraging the spatial redundancy in visual tokens, we proposed a plug-and-play token reduction module that employs the similarity between the class token and spatial tokens as a key criterion for pruning and merging visual tokens.
Our approach, applied to LLaVA-1.5, demonstrated that by utilizing only 6.9% of visual tokens on average, the pruned tokens can maintain comparable performance across a wide range of visual question-answering and reasoning tasks. Notably, our work highlights the potential for significant computational savings without sacrificing the reasoning capabilities of LMMs. We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
6 Limitation and Future Work
Our exploration of LLaVA-PruMerge has two primary limitations. First, the visual token compression, while efficient, is not entirely lossless. This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge. We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps. Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings. This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2023a] framework with a LLaMA-2-34B backbone. Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
Acknowledgement
This work was supported in part by NSF CAREER IIS2150012, and Institute of Information & communications Technology Planning & Evaluation(IITP) grants funded by the Korea government(MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No. RS2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training), and Microsoft Accelerate Foundation Models Research Program.

### v2

@@
 We then cluster the pruned tokens based on key similarity and merge the clustered tokens with the unpruned tokens to supplement their information.
-Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14.4 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
-Code and checkpoints will be released.
+Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 18 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
 1 Introduction
@@
 Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
-Refer to caption (a) Main idea of LLaVA-PruMerge. Refer to caption (b) PruMerged Token Visualization.
+(a) Main idea of LLaVA-PruMerge.
+(b) PruMerged Token Visualization.
 Figure 1: (a) We prune and merge the visual tokens coming from the vision encoder, while keeping all other procedures of the LMM the same.
@@
 Finally, we optionally finetune the LLM to let the model better adapt to our token deduction design.
-Empirically, LLaVA-PruMerge can effectively and adaptively prune the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 6.9% of visual tokens, which is around 40 tokens on average, our model can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
-Our work demonstrates the effectiveness of building efficient large multimodal models from the perspective of visual token pruning and paves the road for further research.
+Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
+Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction and paves the road for further research.
 2 Related Work
-2.1 Large Multimodal Models (LMMs)
+2.1 Efficient Large Multimodal Models (LMMs)
 Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text.
@@
 In this paper, we reduce the number of visual tokens by leveraging the similarity between the class token and the spatial patch tokens.
-2.2 Efficient LMMs
 The need for cross-modal capabilities in resource-limited scenarios has become increasingly important.
@@
 MoE-LLaVA [Lin et al., 2024] adapts Mixture of Experts (MoE) to mitigate model degradation due to sparsity.
+In most cases, LMM efficiency is enhanced by reducing the size of the backbone of the LMM, but no work has considered the efficiency of the LMM from the perspective of the number of visual tokens.
 2.2 Token Reduction
@@
 3.1 Preliminaries
+Figure 2: The conceptual idea of LLaVA-PruMerge.
+Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens; (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster.
+Here $m$ denotes the visual token compression ratio.
 Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024].
@@
 The general forward pass of ViT is illustrated in the left part of Figure 2.
-Figure 2: The conceptual idea of LLaVA-PruMerge.
-Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens; (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster.
-Here $m$ denotes the visual token compression ratio.
 Large Multimodal Models (LMMs).
@@
 3.2 Adaptive Important Token Selection via Outlier Detection
-Figure 3: The distribution of attention values between the [cls] token and visual tokens.
+Figure 3: The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
 The values in the y-axis are logarithmic values.
@@
 Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context.
+Note that we use the class attention value from the penultimate layer for this calculation.
 As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image.
@@
 Such visualizations align with our visual token sampling design.
+These trends are also observed at the benchmark level.
+In Table 3, the average token numbers on various benchmarks differ.
 Algorithm 1 Token PruMerge algorithm for reducing the number of visual tokens.
@@
 This procedure is outlined in Algorithm 1.
+3.4 Discussion
+Distinction from Existing Token Reduction Methods.
+Token reduction methods are especially proposed for inferring vision transformers (ViT) efficiently.
+Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT.
+However, our approach work is not dedigned for efficient ViT, yet it serves as a plug-and-play module after the final outputs of visual encoder (i.e., ViT).
+There are two benefits of this design.
+Firstly, while ViTs typically rely on a single class token to represent the input image, enabling them to maintain performance despite a reduction in intermediate tokens, LMMs usually leverage a large stack of visual tokens.
+This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details.
+Secondly, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not on the reduction of tokens but on maximizing the informational content of the pruned visual tokens.
+This strategy addresses the computational challenges inherent in LMMs without compromising the quality of the visual representation.
 4 Experiments
@@
 Table 1: Comparison with large multimodal models on six benchmarks.
-Our proposed approach LLaVA-PruMerge can adaptively prune and merge visual tokens, which uses only 6.9% visual tokens on average and achieves competitive performance compared to the original LLaVA-1.5 model.
+Our proposed approach LLaVA-PruMerge can adaptively reduce visual tokens, which uses only 5.5% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
 Method LLM Res. PT IT VQAv2 SQAI VQAT POPE MME MMB
@@
 As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 while performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023].
-Specifically, in POPE and SQA, our approach even shows better performance than LLaVA-1.5.
+Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5.
+Note that due to the adaptive nature of PruMerge (see Sec.
+3.2), the token numbers for various tasks are different (see 4.3.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
 4.2 Efficiency Analysis
-To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis as developed in [Yuan et al., 2024].
-Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs.
-Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens.
-Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis.
-Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times to match the token count of the text prompt ($576/14.4\approx 40$).
-The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process.
-Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
 Table 2: Computation Cost Analysis. The development device is TESLA V100 GPU, and time estimated by the roofline model represents the theoretical performance that the hardware can achieve.
@@
 LLaVA-1.5 w/ PruMerge Vicuna-13B INT4 0.45 29.0 6.8 0.11
+To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis as developed in [Yuan et al., 2024].
+Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs.
+Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens.
+Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis.
+Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times in MME/TextVQA to match the token count of the text prompt ($576/14.4\approx 40$).
+The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process.
+Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
 It is important to emphasize that the benefits of PruMerge  extend beyond mere efficiency gains.
@@
 4.3.1 Token Sampling Strategy Analysis
+Figure 4: The demo of the different visual tokens sampling strategies.
+Table 3: Token Sampling Strategy Analysis on Different Tasks.
+Due to the fact that PruMerge can choose the number of visual tokens based on the input image information density, the number of tokens used in different tasks may differ.
+Approach #Visual Tokens Performance
+Task: VQAT
+LLaVA-PruMerge 40 54.00
+Sequential 40 42.72
+Spatial $6\times 6=36$ 46.84
+$5\times 8=40$ 46.85
+$8\times 5=40$ 47.42
+$6\times 7=42$ 47.96
+$7\times 6=42$ 47.49
+Task: MME
+LLaVA-PruMerge 40 1250.07
+Sequential 40 703.60
+Spatial $6\times 6=36$ 1169.10
+$5\times 8=40$ 1180.23
+$8\times 5=40$ 1142.32
+$6\times 7=42$ 1220.89
+$7\times 6=42$ 1199.10
+Task: POPE
+LLaVA-PruMerge 35 76.2
+Sequential 35 11.7
+Spatial $5\times 7=35$ 69.8
+$7\times 5=35$ 71.1
+$6\times 6=36$ 67.9
+Task: SQAI
+LLaVA-PruMerge 16 68.07
+Sequential 16 64.20
+Spatial $4\times 4=16$ 66.29
 Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
-LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec. 3.2), which results in 40 visual tokens per image on average. The visualization is shown in Figure 4 (b).
-Sequential sampling: We sample $N$ tokens in the flatted visual tokens, shown in Figure 4 (c). Here the first 40 tokens are sampled for an apple-to-apple comparison.
-Spatial sampling: Sampled $N$ tokens are spatially distributed evenly across the image, as shown in Figure 4 (d-h).
-We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (36 tokens), and 7 $\times$ 6 (42 tokens).
-Figure 4: The visualization of the different visual tokens sampling strategy.
+LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec.
+3.2), which results in 40 visual tokens per image on average for TextVQA/MME, 35 tokens for POPE, and 16 tokens for SQA.
+The visualization is shown in Figure 4 (b).
+Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apple-to-apple comparison, shown in Figure 4 (c).
+Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h).
+We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens), .
 Note that all the experiments are done via the training-free manner.
 As shown in Table 3, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge.
-Besides, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019].
+Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019].
 Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details.
 This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(a) (b), where more attentive tokens are distributed on the foreground text in the images.
-Table 3: Ablation on training free and fine-tuning for our approach. With further fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
-Approach Number of Visual Tokens SQA TextVQA POPE MME
-Sequential 40 64.60 42.72 13.2 703.60
-Spatial $6\times 6=36$ 66.58 46.84 67.9 1169.10
-$5\times 8=40$ 66.19 46.85 70.4 1180.23
-$8\times 5=40$ 68.12 47.42 71.1 1142.32
-$6\times 7=42$ 67.08 47.96 72.5 1220.89
-$7\times 6=42$ 66.53 47.49 71.6 1199.10
-LLaVA-PruMerge 40 68.07 54.00 76.2 1250.07
 4.3.2 Effectiveness of Each Module in PruMerge
@@
 We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
-6 Limitation and Future Work
+6 Limitations and Future Work
 Our exploration of LLaVA-PruMerge has two primary limitations.
@@
 Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings.
-This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2023a] framework with a LLaMA-2-34B backbone.
+This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2023a] framework with a Yi-34B backbone.
 Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.

### v3

@@
 In Table 3, the average token numbers on various benchmarks differ.
-Algorithm 1 Token PruMerge algorithm for reducing the number of visual tokens.
+Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
 1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$.
@@
 5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ index $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
-6:for $p=\{i_{1},\cdots,i_{m}\}$ do
-7:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
-8:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
-9:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
-10:end for
-11:(see Sec. 3.3).
-12:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
+6:(Optional for PruMerge+) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
+7:(Optional for PruMerge+) Spacial-uniformly sample visual tokens with $r_{o}=\frac{m}{n}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
+8:(Optional for PruMerge+) Update the selected tokens index with $\{i_{1},\cdots,i_{m},1_{1+1},\cdots,i_{2m}\}$.
+9:for $p=\{i_{1},\cdots,i_{m}\}$ do
+10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
+11:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
+12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
+13:end for
+14:(see Sec. 3.3).
+15:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
 3.3 Token Supplement via Similar Key Clustering
@@
 This procedure is outlined in Algorithm 1.
+3.4 PruMerge+: Bridging the Efficiency-Performance Gap
+While PruMergeachieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without its drawbacks.
+Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge.
+To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
+Figure 4: Image Visualizations of PruMerge and PruMerge+.
+PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation.
+This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2.
+A new aspect of PruMerge+ lies in its enhanced token selection process.
+Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical.
+This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens.
+This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 4, thereby minimizing any performance losses while still achieving substantial token reduction.
 3.5 Discussion
@@
 LLaVA-1.5 Vicuna-7B 336 558K 665K 78.5 66.8 58.2 85.9 1510.7 64.3
-LLaVA-1.5 + Ours Vicuna-7B 336 558K 665K 72.0 68.5 56.0 86.3 1350.3 60.9
+LLaVA-1.5 + PruMerge Vicuna-7B 336 558K 665K 72.0 68.5 56.0 76.3 1350.3 60.9
+LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.5 68.3 57.1 84.0 1431.6 63.5
 LLaVA-1.5 Vicuna-13B 336 558K 665K 80.0 71.6 61.3 85.9 1531.3 67.7
-LLaVA-1.5 + Ours Vicuna-13B 336 558K 665K 72.8 71.0 58.4 86.2 1428.2 62.3
+LLaVA-1.5 + PruMerge Vicuna-13B 336 558K 665K 72.8 71.0 58.4 78.5 1428.2 62.3
+LLaVA-1.5 + PruMerge+ Vicuna-13B 336 558K 665K 77.8 71.0 58.6 84.4 1485.5 65.7
 4.1 Main Results
@@
 Figure 5: The demo of the different visual tokens sampling strategies.
+Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
 Table 3: Token Sampling Strategy Analysis on Different Tasks.
@@
 Spatial $4\times 4=16$ 66.29
-Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
 LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec.
@@
 LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
-LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 85.7 1221.6
-LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 86.3 1350.3
+LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
+LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
 Here, we study the effectiveness of each module in our design based on LLaVA-1.5.
@@
 LLaVA-PruMerge w.o. fine-tuning Vicuna-7B 68.0 54.0 76.2 1250.1
-LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 86.3 1350.3
+LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 76.3 1350.3
 5 Conclusion
@@
 We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
-6 Limitations and Future Work
-Our exploration of LLaVA-PruMerge has two primary limitations.
-First, the visual token compression, while efficient, is not entirely lossless.
-This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge.
-We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps.
-Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings.
-This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2023a] framework with a Yi-34B backbone.
-Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
 Acknowledgement

### v4

@@
 Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
-Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction and paves the road for further research.
+Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction.
 2 Related Work
@@
 In this paper, we reduce the number of visual tokens by leveraging the similarity between the class token and the spatial patch tokens.
-The need for cross-modal capabilities in resource-limited scenarios has become increasingly important.
-Despite advancements in LMMs, their large-scale training and deployment incur significant computational costs, necessitating efficient parallel device implementations.
-Google’s Gemini [Team et al., 2023] is a leader in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices.
-However, Gemini remains closed-source.
-Open-source initiatives, like LLaVA-1.5 [Liu et al., 2023a], utilize advanced compression techniques, such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024].
-Further efforts towards efficient LMMs include MobileVLM [Chu et al., 2023], which develops a compact LLM and an efficient multimodal feature projector, and its successor, MobileVLM-v2 [Chu et al., 2024], which explores improved training strategies for mobile scenarios.
-TinyGPT-V [Yuan et al., 2023a] leverages the advanced Phi-2 [Javaheripi et al., 2023] LLM to surpass the performance of significantly larger models.
-Similarly, LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for broader generalizability.
-TinyLLaVA [Zhou et al., 2024] investigates the impacts of architectural choices, data quality, and training strategies, demonstrating that smaller LMMs can match the performance of their larger counterparts with optimized data and training.
-MoE-LLaVA [Lin et al., 2024] adapts Mixture of Experts (MoE) to mitigate model degradation due to sparsity.
+While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations.
+Google’s Gemini [Team et al., 2023] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source.
+Open-source alternatives like LLaVA-1.5 [Liu et al., 2023a] employ advanced compression techniques such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024].
+MobileVLM [Chu et al., 2023] and its improved version, MobileVLM-v2 [Chu et al., 2024], focus on compact architecture designs and training optimizations for mobile use.
+Similarly, TinyGPT-V [Yuan et al., 2023a] uses the Phi-2 [Javaheripi et al., 2023] LLM backbone to exceed the performance of larger models.
+LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for better generalizability.
+TinyLLaVA [Zhou et al., 2024] explores the effects of architectural choices and training optimizations, achieving comparable performance to larger models.
+MoE-LLaVA [Lin et al., 2024] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance.
 In most cases, LMM efficiency is enhanced by reducing the size of the backbone of the LMM, but no work has considered the efficiency of the LMM from the perspective of the number of visual tokens.
@@
 Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024].
-ViTs perform tokenization by dividing an input image into patches and projecting each patch into a token embedding.
-An extra class token [CLS] is added to the set of image tokens and is responsible for aggregating global image information and performing classification.
+An input image is divided into patches and each patch is projected into a token embedding by ViTs.
+In addition to the set of image tokens, a new class token (i.e., [CLS] token) is added to aggregate global image information and perform classification.
 Let $n$ denote the number of input tokens (image patches) to the ViT.
-ViTs are composed of a series of transformer blocks.
-An assembly of several key components composes each transformer block: a multi-head self-attention layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016], all of which work together to enhance the model’s ability to capture and interpret visual information [Han et al., 2022].
-In the self-attention layer, an input token is projected into three distinct vectors: the query vector $\mathbf{q}$, the key vector $\mathbf{k}$, and the value vector $\mathbf{v}$, utilizing three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$.
+Let $n$ denote the number of input tokens to the ViT.
+ViTs consist of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016].
+These components work together to improve the model’s capability to understand visual data [Han et al., 2022].
+In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$.
 These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$.
-The self-attention mechanism computes the relevance of each item to other items as follows:
+The self-attention computes the relevance of each item to other items:
 $\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (3.1)
-where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ represents the dimension of $\mathbf{q},\mathbf{k},\mathbf{v}$.
-In the last encoder layer of ViT, the [CLS] token is used for classification.
-Similarly, the attention between [CLS] token and other visual tokens are performed via the attention mechanism:
+where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$.
+In the last layer of ViT, the [CLS] token is used for classification.
+Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
 $\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (3.2)
-The multi-head self-attention framework allows for simultaneous focus on multiple positions, offering diverse representation subspaces.
-This is achieved by employing distinct query, key, and value matrices for different heads, which project the input vectors into varied representation subspaces.
-Subsequent to the self-attention layers is the feed-forward network (FFN).
-This network consists of two linear transformation layers separated by a nonlinear activation function and is represented as:
+The MSA framework allows for simultaneous attention on multiple positions, offering diverse representation subspaces.
+This is achieved by employing distinct query, key, and value matrices for different heads, which project the input vectors into different representation subspaces.
+After the self-attention layers is the feed-forward network (FFN), which consists of two linear transformation layers separated by a nonlinear activation function:
 $\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3.3)
-with $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ being the parameter matrices of the linear transformation layers and $\sigma$ denoting the nonlinear activation function, such as GELU [Hendrycks & Gimpel, 2016].
+where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, as well as $\sigma$ denoting the nonlinear activation function.
 The general forward pass of ViT is illustrated in the left part of Figure 2.
 Large Multimodal Models (LMMs).
-Following the forward pass through a Vision Transformer (ViT), a series of visual tokens is generated.
-These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which is responsible for mapping the encoded features from other modalities $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$.
+Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated.
+These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which can map the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$.
 The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024].
 The overall architecture of LMMs is depicted in Figure 1.
-It is important to note that the computational cost associated with these models tends to increase quadratically with the number of input tokens fed into the LLM backbone [Tay et al., 2022].
-There is an increasing demand for the processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs.
+It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022].
+There is an increasing demand for processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs.
 The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
 3.2 Adaptive Important Token Selection via Outlier Detection
-Figure 3: The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
+(a) Distribution of attention values.
+(b) Image & Tokens Visualizations
+Figure 3: (a) The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
 The values in the y-axis are logarithmic values.
 Notably, most of the spatial visual tokens have near-zero attention values with the class token.
+(b) Image Visualizations of PruMerge and PruMerge+.
 The most straight-forward solution to improve the efficiency of visual token utilization in LMMs is to prune the redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022].
@@
 To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2.
-Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3.
+Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3(a).
 This sparse distribution underpins our methodology for identifying crucial visual tokens.
@@
 To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens.
-Luckily, transformers natively solve this problem with QKV self-attention.
-Specifically, the keys (K) already summarize the information contained in each token for use in dot product similarity.
-Thus, we use a dot product similarity metric (e.g., cosine similarity) between the keys of each token to determine which tokens contain similar visual information [Bolya et al., 2023].
+Here we leverage the self-attention mechanism in Vision Transformers (ViT) to metric such a similarity.
+Specifically, the final layer’s key vector is utilized as the representation of each image token, and thus we use a dot product similarity metric of keys to calculate which tokens have similar visual information [Bolya et al., 2023].
 Specifically, the similarity between visual tokens is computed as
@@
 To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
-Figure 4: Image Visualizations of PruMerge and PruMerge+.
 PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation.
@@
 This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens.
-This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 4, thereby minimizing any performance losses while still achieving substantial token reduction.
+This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing any performance losses while still achieving substantial token reduction.
 3.5 Discussion
@@
 Table 1: Comparison with large multimodal models on six benchmarks.
-Our proposed approach LLaVA-PruMerge can adaptively reduce visual tokens, which uses only 5.5% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
+Our proposed approach PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
 Method LLM Res. PT IT VQAv2 SQAI VQAT POPE MME MMB
@@
 LLaVA-1.5 + PruMerge Vicuna-7B 336 558K 665K 72.0 68.5 56.0 76.3 1350.3 60.9
-LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.5 68.3 57.1 84.0 1431.6 63.5
+LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.8 68.3 57.1 84.0 1462.4 64.9
 LLaVA-1.5 Vicuna-13B 336 558K 665K 80.0 71.6 61.3 85.9 1531.3 67.7
@@
 Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
+LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec.
+3.2), which results in 40 visual tokens per image on average for TextVQA/MME, 35 tokens for POPE, and 16 tokens for SQA.
+The visualization is shown in Figure 4 (b).
+Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apple-to-apple comparison, shown in Figure 4 (c).
+Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h).
+We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
 Table 3: Token Sampling Strategy Analysis on Different Tasks.
@@
 Spatial $4\times 4=16$ 66.29
-LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec.
-3.2), which results in 40 visual tokens per image on average for TextVQA/MME, 35 tokens for POPE, and 16 tokens for SQA.
-The visualization is shown in Figure 5 (b).
-Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apple-to-apple comparison, shown in Figure 5 (c).
-Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 5 (d-h).
-We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens), .
+Table 4: Ablation Studies for Adaptive Important Token Selection (AITS, Sec. 3.2) and Token Supplement (TS, Sec. 3.3). With these modules, the downstream performance can be progressively improved.
+Method LLM SQAI VQAT POPE MME
+LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
+LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
+LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
 Note that all the experiments are done via the training-free manner.
@@
 4.3.2 Effectiveness of Each Module in PruMerge
-Table 4: Ablation Studies for Adaptive Important Token Selection (AITS, Sec. 3.2) and Token Supplement (TS, Sec. 3.3). With these modules, the downstream performance can be progressively improved.
-Method LLM SQAI VQAT POPE MME
-LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
-LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
-LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
 Here, we study the effectiveness of each module in our design based on LLaVA-1.5.

### v5

@@
 Abstract
-Large Multimodal Models (LMMs) have shown significant reasoning capabilities by connecting a visual encoder and a large language model.
-LMMs typically use a fixed amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content.
-Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which increase the number of visual tokens significantly.
-However, due to the design of the Transformer architecture, computational costs associated with these models tend to increase quadratically with the number of input tokens.
-To tackle this problem, we explore a token reduction mechanism and find, similar to prior work, that many visual tokens are spatially redundant.
-Based on this, we propose PruMerge, a novel adaptive visual token reduction approach, which largely reduces the number of visual tokens while maintaining comparable model performance.
-We first select the unpruned visual tokens based on their similarity to class tokens and spatial tokens.
-We then cluster the pruned tokens based on key similarity and merge the clustered tokens with the unpruned tokens to supplement their information.
-Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 18 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
+Large Multimodal Models (LMMs) have shown significant visual reasoning capabilities by connecting a visual encoder and a large language model.
+LMMs typically take in a fixed and large amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content.
+Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which further increases the number of visual tokens significantly.
+However, due to the inherent design of the Transformer architecture, the computational costs of these models tend to increase quadratically with the number of input tokens.
+To tackle this problem, we explore a token reduction mechanism that identifies significant spatial redundancy among visual tokens.
+In response, we propose PruMerge, a novel adaptive visual token reduction strategy that significantly reduces the number of visual tokens without compromising the performance of LMMs.
+Specifically, to metric the importance of each token, we exploit the sparsity observed in the visual encoder, characterized by the sparse distribution of attention scores between the class token and visual tokens.
+This sparsity enables us to dynamically select the most crucial visual tokens to retain.
+Subsequently, we cluster the selected (unpruned) tokens based on their key similarity and merge them with the unpruned tokens, effectively supplementing and enhancing their informational content.
+Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
 1 Introduction
 Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities.
-LLMs are usually high-capacity Transformers pretrained with a large-scale text corpus.
-Large Multimodal Models (LMMs), inherit the LLMs for text generation, but also leverage a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
-LMMs need substantial computation to conduct inference.
+LLMs are usually high-capacity Transformers [Vaswani et al., 2017] pretrained with a large-scale text corpus.
+Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
+LMMs need substantial computation for inference.
 The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM.
@@
 Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023].
-However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap in visual question-answering and reasoning such as VQAv2 and MM-Bench [Chu et al., 2024].
+However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [Chu et al., 2024].
 A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
@@
 The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens.
-For example, in LLaVA-1.5, 576 visual tokens are appended, leading to high training and inference costs.
+For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [Lin et al., 2023] that number is even higher, leading to high training and inference costs.
 Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
-(a) Main idea of LLaVA-PruMerge.
+(a) Main idea of PurMerge.
 (b) PruMerged Token Visualization.
-Figure 1: (a) We prune and merge the visual tokens coming from the vision encoder, while keeping all other procedures of the LMM the same.
-By reducing the number of visual tokens, our proposed method, PruMerge, significantly reduces the computation cost for text generation in LMMs.
-(b) The visualizations of the attentive tokens.
-We design a token reduction method to adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently.
+Figure 1: (a) We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same.
+By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance.
+(b) A visualization of the selected tokens.
+PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently.
 More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images.
-Besides, such attentive tokens are usually located at the regions with dense information.
-In our study, we find that such visual tokens are redundant, similar to findings in previous related work Bolya et al.
-[2023], Liu et al.
-[2022], and most of the visual tokens can be pruned without largely sacrificing the performance.
-In particular, we find that the activations are very sparse upon the similarity matrix between the class token and spatial patches, which indicates that only a small amount of the visual tokens are related to the key visual information in the image.
-Motivated by this, we use this similarity to select important visual tokens.
-Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune the visual tokens.
-Moreover, we merge the visual tokens using $k$-nearest neighbor and update the sampled visual tokens via weighted averaging, which further enhances performance.
-Finally, we optionally finetune the LLM to let the model better adapt to our token deduction design.
-Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
-Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction.
+The attentive tokens are usually located at the regions with dense information.
+In our study, we find that many visual tokens are redundant, similar to findings in previous related work [Bolya et al., 2023, Liu et al., 2022], and most of the visual tokens can be pruned with little sacrifice in performance.
+In particular, the similarity (i.e., attention scores in the visual encoder’s self-attention module) between the class token and spatial patches are sparse, indicating that only a small number of visual tokens are related to key visual information in the visual samples.
+Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1(b).
+Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune unimportant visual tokens.
+Moreover, we merge the visual tokens using $k$-nearest neighbor and update the selected important visual tokens via weighted averaging, which further enhances performance.
+Finally, we design PruMerge+, which samples visual tokens spatial-uniformly to complement the unpruned tokens.
+PruMerge+ not only minimizes performance degradation but also ensures substantial token reduction, maintaining a more comprehensive and representative selection of visual tokens.
+Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
+Furthermore, PruMerge showcases its versatility across various modalities, including video.
+By integrating PruMerge with Video-LLaVA during the inference phase alone—eliminating the need for additional training—we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
 2 Related Work
@@
 More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023].
-Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022].
+Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022].
 The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a].
-In this paper, we reduce the number of visual tokens by leveraging the similarity between the class token and the spatial patch tokens.
+In this paper, we reduce the number of visual tokens with a novel adaptive prune and merge procedure.
 While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations.
@@
 2.2 Token Reduction
-The notorious squared complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the sequence length.
-Sparse attention such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context.
+The notorious quadratic complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the input sequence length.
+Sparse attention methods such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context.
 Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching.
-In recent LMMs [Liu et al., 2023b, Zhu et al., 2023], prefix visual tokens serve as a fixed budget for context, which becomes one of the leading factors for their efficiency.
-In our study, we find that by simply pruning and merging visual tokens based on their similarity, we can achieve comparable performance using less than one tenth of the original tokens.
+One of the main causes of the inefficiency of recent LMMs is their use of a large number of prefix visual tokens that serve as a fixed budget for context [Liu et al., 2023b, Zhu et al., 2023].
+In our study, we present a novel plug-and-play token reduction method based on visual token similarities, which achieves comparable performance with less than one tenth of the original tokens.
 3 Method: Token Pru-Merging
@@
 Our method features two key components: (1) Adaptive Important Token Selection (AITS) via Outlier Detection which adaptively determines the optimal number of visual tokens to retain based on the unique characteristics of the image (Sec.
-3.2); and (2) Token Supplement (TS) via Similar Key Clustering, which facilitates more efficient processing without compromising the model’s performance by maintaining the integrity and richness of the visual information (Sec.
+3.2); and (2) Token Supplement (TS) via Similar Key Clustering, which facilitates efficient processing without compromising the model’s performance by maintaining the integrity and richness of the visual information (Sec.
 3.3).
 3.1 Preliminaries
-Figure 2: The conceptual idea of LLaVA-PruMerge.
-Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens; (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster.
+Figure 2: The conceptual idea of PruMerge.
+Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3).
 Here $m$ denotes the visual token compression ratio.
-Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024].
-An input image is divided into patches and each patch is projected into a token embedding by ViTs.
-In addition to the set of image tokens, a new class token (i.e., [CLS] token) is added to aggregate global image information and perform classification.
-Let $n$ denote the number of input tokens (image patches) to the ViT.
-Let $n$ denote the number of input tokens to the ViT.
-ViTs consist of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016].
+Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024].
+An input image is divided into a grid of patches and each patch is projected into a token embedding by the ViT.
+In addition to the patch tokens, a class token (i.e., [CLS] token) is computed to aggregate global image information for classification.
+A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [Ba et al., 2016].
 These components work together to improve the model’s capability to understand visual data [Han et al., 2022].
 In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$.
-These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$.
+These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, respectively.
 The self-attention computes the relevance of each item to other items:
 $\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (3.1)
-where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$.
-In the last layer of ViT, the [CLS] token is used for classification.
+where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$ and $\mathbf{k}$.
+In the last layer of the ViT, the [CLS] token is used for classification.
 Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
@@
 $\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3.3)
-where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, as well as $\sigma$ denoting the nonlinear activation function.
+where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, and $\sigma$ denotes the nonlinear activation function.
 The general forward pass of ViT is illustrated in the left part of Figure 2.
@@
 Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated.
-These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which can map the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$.
+These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which maps the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$.
 The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024].
-The overall architecture of LMMs is depicted in Figure 1.
+The overall architecture of an LMM is depicted in Figure 1.
 It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022].
-There is an increasing demand for processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs.
+Mathematically, if there are $N$ tokens in the input, the self-attention mechanism computes a $N\times N$ matrix of attention scores, where each entry in this matrix represents the attention score between a pair of tokens.
+However, there is an increasing demand for processing high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs.
 The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@
 (b) Image & Tokens Visualizations
-Figure 3: (a) The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
+Figure 3: (a) The distribution of attention scores (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
 The values in the y-axis are logarithmic values.
 Notably, most of the spatial visual tokens have near-zero attention values with the class token.
-(b) Image Visualizations of PruMerge and PruMerge+.
-The most straight-forward solution to improve the efficiency of visual token utilization in LMMs is to prune the redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022].
+(b) Visualizations of PruMerge and PruMerge+.
+The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022].
 To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
@@
 Any attention values residing outside these fences are classified as outliers.
+In practice, only the “upper fence” is activated since the attention score is positive.
 Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context.
@@
 As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image.
-Simpler images such as ”Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens.
-(2) The sampled tokens are spatially aligned with the important details, while less tokens are assigned for the background.
+Simpler images such as “Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens.
+(2) The sampled tokens are typically spatially aligned with important content.
 Such visualizations align with our visual token sampling design.
-These trends are also observed at the benchmark level.
-In Table 4, the average token numbers on various benchmarks differ.
-Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
-1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$.
-The penultimate layer’s output tokens, $\mathbf{Y}=\{\mathbf{y}_{1},\cdots\mathbf{y}_{n}\}$.
-$n$ is the number of input visual tokens.
-2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
-3:Token PruMerge:
-4:Calculate attention between visual token and class token $\mathbf{a}_{[cls]}$ using Equation 3.2.
-5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ index $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
-6:(Optional for PruMerge+) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
-7:(Optional for PruMerge+) Spacial-uniformly sample visual tokens with $r_{o}=\frac{m}{n}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
-8:(Optional for PruMerge+) Update the selected tokens index with $\{i_{1},\cdots,i_{m},1_{1+1},\cdots,i_{2m}\}$.
-9:for $p=\{i_{1},\cdots,i_{m}\}$ do
-10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
-11:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
-12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
-13:end for
-14:(see Sec. 3.3).
-15:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
+These trends are also observed at the benchmark level; in Table 6, the average token numbers on various benchmarks differ.
 3.3 Token Supplement via Similar Key Clustering
-Following the selection of more informative visual tokens, we next optimize the utilization of the remaining tokens.
+Following the selection of informative visual tokens, we next optimize the utilization of the remaining tokens.
 While pruned tokens may initially seem extraneous, they hold potential value for the perception capabilities of the LLM backbone.
@@
 To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens.
-Here we leverage the self-attention mechanism in Vision Transformers (ViT) to metric such a similarity.
-Specifically, the final layer’s key vector is utilized as the representation of each image token, and thus we use a dot product similarity metric of keys to calculate which tokens have similar visual information [Bolya et al., 2023].
-Specifically, the similarity between visual tokens is computed as
+Here we leverage the self-attention mechanism in ViTs.
+Since the key vector of each patch token already contains information summarized in the self-attention module [Vaswani et al., 2017], the final layer’s key vector serves as the representation.
+And then we use the dot product between keys to calculate which tokens have similar visual information [Bolya et al., 2023]:
 $\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (3.4)
-which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorization form for the set of all tokens ${1,2,\cdots,n}$.
+which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorized form for the set of all tokens ${1,2,\cdots,n}$, where $n$ is the number of input visual tokens.
 With the similarities between visual tokens established, we simply find the $k$-nearest neighbors for each unpruned token, which act as the cluster centers.
@@
 3.4 PruMerge+: Bridging the Efficiency-Performance Gap
-While PruMergeachieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without its drawbacks.
-Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge.
+While PruMerge achieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without drawbacks.
+Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the original LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge.
 To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
-PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation.
+PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—with minimal performance degradation.
 This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2.
@@
 This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens.
-This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing any performance losses while still achieving substantial token reduction.
+This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing performance losses while still achieving substantial token reduction.
+Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
+1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$.
+The penultimate layer’s output tokens, $\mathbf{Y}=\{\mathbf{y}_{1},\cdots\mathbf{y}_{n}\}$.
+$n$ is the number of input visual tokens.
+2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
+3:Token PruMerge:
+4:Calculate attention between visual token and class token $\mathbf{a}_{[cls]}$ using Equation 3.2.
+5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ indices $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
+6:(Optional for PruMerge+, see Sec. 3.4) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
+7:(Optional for PruMerge+) Spatial-uniformly sample visual tokens with $r_{o}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
+8:(Optional for PruMerge+) Update the selected tokens’ index with $\{i_{1},\cdots,i_{m},i_{m+1},\cdots,i_{2m}\}$.
+9:for $p=\{i_{1},\cdots,i_{m}\}$ do (see Sec. 3.3)
+10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
+11:     Use $\mathbf{y}_{p}$ as cluster center and find the $k$ most similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
+12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
+13:end for
+14:Output a refined stack of visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
 3.5 Discussion
 Distinction from Existing Token Reduction Methods.
-Token reduction methods are especially proposed for inferring vision transformers (ViT) efficiently.
+Token reduction methods [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022] have been proposed for accelerating ViT computation speed.
 Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT.
-However, our approach work is not dedigned for efficient ViT, yet it serves as a plug-and-play module after the final outputs of visual encoder (i.e., ViT).
+By contrast, our approach is not specifically designed to increase the efficiency of a ViT.
+It instead aims to improve the overall efficiency of a large multimodal model, where ViT is only one relatively light cost component as shown in Fig.
+1.
 There are two benefits of this design.
-Firstly, while ViTs typically rely on a single class token to represent the input image, enabling them to maintain performance despite a reduction in intermediate tokens, LMMs usually leverage a large stack of visual tokens.
+First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens.
 This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details.
-Secondly, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not on the reduction of tokens but on maximizing the informational content of the pruned visual tokens.
-This strategy addresses the computational challenges inherent in LMMs without compromising the quality of the visual representation.
+Thus, using previous token merging methods to obtain one refined class token as representation of visual input is not consistent with the literature of large multimodal models.
+Second, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens.
+This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation.
 4 Experiments
-We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then demonstrate the effectiveness of each component in our model in Sec 4.3.
+We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1.
+We then analyze the efficiency improvement by using our PruMerge on LMM in Sec 4.2.
+To show the generalization ablity, we conduct a series of experiments in Sec.
+4.3.
+Finally, we demonstrate the effectiveness of each component in our model in Sec 4.4.
 Table 1: Comparison with large multimodal models on six benchmarks.
@@
 We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c].
-As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 while performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023].
+As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023].
 Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5.
 Note that due to the adaptive nature of PruMerge (see Sec.
-3.2), the token numbers for various tasks are different (see 4.3.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
+3.2), the token numbers for various tasks are different (see 4.4.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
 4.2 Efficiency Analysis
@@
 Method LLM Quantization FLOPs Prefill Total Storing
-Backbone (T) Time (ms) Memory (G) Activation (G)
+Backbone (TB) Time (ms) Memory (GB) Activation (GB)
 LLaVA-1.5 Vicuna-7B FP16 9.3 88.6 23.3 4.60
@@
 LLaVA-1.5 w/ PruMerge Vicuna-13B INT4 0.45 29.0 6.8 0.11
-To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis as developed in [Yuan et al., 2024].
+To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [Yuan et al., 2024].
 Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs.
@@
 This orthogonal relationship underscores the versatile potential of PruMerge to contribute to a broader spectrum of efficiency-enhancing strategies.
+4.3 Generalization on Video-LLM
+To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [Lin et al., 2023].
+Video-LLaVA is one of the most popular open-soruced Video-LLMs.
+We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference.
+The outcome of this integration is shown in Table 3.
+Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [Liu et al., 2023a].
+Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively.
+The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance.
+This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs.
+Exploring ways to further capitalize on this redundancy could shape future research directions.
+Table 3: A comparison of different LVMs on video reasoning benchmarks.
+Like Video-LLaVA, ChatGPT-Assistant is employed to evaluate performance.
+The version of ChatGPT is “gpt-3.5-turbo”.
+Note that we directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
+Methods LLM size MSVD-QA MSRVT-QA ActivityNet-QA
+Accuracy Score Accuracy Score Accuracy Score
+FrozenBiLM 1B 32.2 - 16.8 - 24.7 -
+VideoChat 7B 56.3 2.8 45.0 2.5 - 2.2
+LLaMA-Adapter 7B 54.9 3.1 43.8 2.7 34.2 2.7
+Video-LLaMA 7B 51.6 2.5 29.6 1.8 12.4 1.1
+Video-ChatGPT 7B 64.9 3.3 49.3 2.8 35.2 2.7
+Video-LLaVA 7B 70.7 3.9 59.2 3.5 45.3 3.3
+Video-LLaVA + PruMerge 7B 71.1 3.9 58.4 3.5 48.3 3.4
+Video-LLaVA + PruMerge+ 7B 71.1 3.9 59.3 3.6 47.7 3.4
 4.4 Ablation Study
 4.4.1 Token Sampling Strategy Analysis
-Figure 4: The demo of the different visual tokens sampling strategies.
 Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
@@
 The visualization is shown in Figure 4 (b).
-Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apple-to-apple comparison, shown in Figure 4 (c).
+Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apples-to-apples comparison, shown in Figure 4 (c).
+Figure 4: Different token sampling strategies.
 Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h).
-We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
+We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
 Table 4: Token Sampling Strategy Analysis on Different Tasks.
-Due to the fact that PruMerge can choose the number of visual tokens based on the input image information density, the number of tokens used in different tasks may differ.
 Approach #Visual Tokens Performance
@@
 Sequential 40 42.72
-Spatial $6\times 6=36$ 46.84
-$5\times 8=40$ 46.85
+Spatial $5\times 8=40$ 46.85
 $8\times 5=40$ 47.42
-$6\times 7=42$ 47.96
-$7\times 6=42$ 47.49
 Task: MME
@@
 Sequential 40 703.60
-Spatial $6\times 6=36$ 1169.10
-$5\times 8=40$ 1180.23
+Spatial $5\times 8=40$ 1180.23
 $8\times 5=40$ 1142.32
-$6\times 7=42$ 1220.89
-$7\times 6=42$ 1199.10
 Task: POPE
@@
 LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
-LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
-LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
-Note that all the experiments are done via the training-free manner.
+LLaVA-1.5 w. AITS Vicuna-7B 66.5 54.8 75.7 1221.6
+LLaVA-1.5 w. AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
+Table 6: Ablation on training free and fine-tuning for our approach. With fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
+Method LLM SQAI VQAT POPE MME
+LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
+LLaVA-PruMerge w.o. LoRA-FT Vicuna-7B 68.0 54.0 76.2 1250.1
+LLaVA-PruMerge w. LoRA-FT Vicuna-7B 68.5 56.0 76.3 1350.3
+Note that all the experiments are done via a training-free manner.
 As shown in Table 6, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge.
@@
 Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details.
-This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(a) (b), where more attentive tokens are distributed on the foreground text in the images.
+This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(b), where more attentive tokens are distributed on the foreground text in the images.
 4.4.2 Effectiveness of Each Module in PruMerge
@@
 4.4.3 Training Analysis: Training-free v.s. Fine-tuning
-LLaVA-PruMerge can be conducted both in training-free and training-needed manners.
+Finally, LLaVA-PruMerge can be conducted in either a training-free or fine-tuning manner.
 With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks.
 As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].
-Table 5: Ablation on training free and fine-tuning for our approach. With further fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
-Method LLM SQAI VQAT POPE MME
-LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
-LLaVA-PruMerge w.o. fine-tuning Vicuna-7B 68.0 54.0 76.2 1250.1
-LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 76.3 1350.3
 5 Conclusion
@@
 We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
-Acknowledgement
-This work was supported in part by NSF CAREER IIS2150012, and Institute of Information & communications Technology Planning & Evaluation(IITP) grants funded by the Korea government(MSIT) (No.
-2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No.
-RS2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training), and Microsoft Accelerate Foundation Models Research Program.
+Appendix A Appendix
+A.1 Limitations and Future Work
+Our exploration of LLaVA-PruMerge has two primary limitations.
+First, the visual token compression, while efficient, is not entirely lossless.
+This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge.
+We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps.
+Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings.
+This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2024] framework with a Yi-34B backbone.
+Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
+A.2 Societal Impacts
+In this study, we propose a technique that improves the efficiency of Large Multimodal Models (LMMs), making them more accessible.
+This approach helps to democratize LMMs by lowering deployment costs and hardware barriers, facilitating their use in edge computing.
+However, it does not mitigate the potential misuse of LMMs by malicious actors.

### v6

@@
 Subsequently, we cluster the selected (unpruned) tokens based on their key similarity and merge them with the unpruned tokens, effectively supplementing and enhancing their informational content.
-Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
+Empirically, when applied to LLaVA-1.5 [28] and Video-LLaVA [26], our approach can reduce the number of visual tokens by 4 times, and achieve comparable or better performance across diverse visual question-answering and reasoning tasks.
 1 Introduction
-Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities.
-LLMs are usually high-capacity Transformers [Vaswani et al., 2017] pretrained with a large-scale text corpus.
-Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
+Large Language Models (LLMs) [34, 42, 21, 43] have shown strong reasoning abilities.
+LLMs are usually high-capacity Transformers [44] pretrained with a large-scale text corpus.
+Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [37] to embed image patches into visual tokens as the prefix visual context.
+Figure 1: Main idea of PruMerge: We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same.
+By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance.
+A visualization of the selected tokens: PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently.
+More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images.
+Attentive tokens are located at the regions with dense information.
 LMMs need substantial computation for inference.
 The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM.
-For example, the commonly used CLIP visual encoder, ViT-L, only has 0.3B parameters, while the corresponding LLM such as LLaMA [Touvron et al., 2023] or Vicuna [Vicuna, 2023] can have 7B or 13B parameters.
-As a result, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
-Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023].
-However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [Chu et al., 2024].
-A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
+For example, the commonly used CLIP visual encoder only has 0.3B parameters, while the corresponding LLM such as LLaMA [43], Vicuna [45], DeepSeek [27], and Qwen [3] can have 7B, 13B, or even 70B+ parameters.
+Therefore, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
+Prior works [8, 9, 49] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [20].
+However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [9].
+A similar approach is to apply quantization for LLMs [30, 51].
 However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation.
 The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens.
-For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [Lin et al., 2023] that number is even higher, leading to high training and inference costs.
+For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [26] that number is even higher (2048), leading to high training and inference costs.
 Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
-(a) Main idea of PurMerge.
-(b) PruMerged Token Visualization.
-Figure 1: (a) We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same.
-By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance.
-(b) A visualization of the selected tokens.
-PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently.
-More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images.
-The attentive tokens are usually located at the regions with dense information.
-In our study, we find that many visual tokens are redundant, similar to findings in previous related work [Bolya et al., 2023, Liu et al., 2022], and most of the visual tokens can be pruned with little sacrifice in performance.
+In our study, we find that many visual tokens are redundant, similar to findings in previous related work [4, 31], and most of the visual tokens can be pruned with little sacrifice in performance.
 In particular, the similarity (i.e., attention scores in the visual encoder’s self-attention module) between the class token and spatial patches are sparse, indicating that only a small number of visual tokens are related to key visual information in the visual samples.
-Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1(b).
-Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune unimportant visual tokens.
+Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1.
+Specifically, we leverage the Interquartile Range (IQR) [5] scoring function in outlier detection to prune unimportant visual tokens.
 Moreover, we merge the visual tokens using $k$-nearest neighbor and update the selected important visual tokens via weighted averaging, which further enhances performance.
@@
 PruMerge+ not only minimizes performance degradation but also ensures substantial token reduction, maintaining a more comprehensive and representative selection of visual tokens.
-Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
+Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [28], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks.
 Furthermore, PruMerge showcases its versatility across various modalities, including video.
-By integrating PruMerge with Video-LLaVA during the inference phase alone—eliminating the need for additional training—we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
+By integrating PruMerge with Video-LLaVA [26] during the inference phase alone (i.e., no need for additional training) we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
 2 Related Work
 2.1 Efficient Large Multimodal Models (LMMs)
-Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text.
-Large Multimodal Models (LMMs) [Liu et al., 2023b, Zhu et al., 2023, Yin et al., 2023, Zhang et al., 2024] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format.
-More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023].
-Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022].
-The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a].
+Large Language Models (LLMs) such as GPT-4 [34], LLaMA [43], Mistral [21], and Gemini [42] have demonstrated strong question answering and reasoning capabilities over text.
+Large Multimodal Models (LMMs) [30, 55, 48, 52] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format.
+More recent works extend whole-image understanding into region-level understanding [6, 54, 36, 7], video understanding [26, 53] and 3D scene understanding [18].
+Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [28], Qformer [10, 55], or resampler [1].
+The number of visual tokens can be prohibitively long, especially when the images are high-resolution [29, 35].
 In this paper, we reduce the number of visual tokens with a novel adaptive prune and merge procedure.
 While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations.
-Google’s Gemini [Team et al., 2023] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source.
-Open-source alternatives like LLaVA-1.5 [Liu et al., 2023a] employ advanced compression techniques such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024].
-MobileVLM [Chu et al., 2023] and its improved version, MobileVLM-v2 [Chu et al., 2024], focus on compact architecture designs and training optimizations for mobile use.
-Similarly, TinyGPT-V [Yuan et al., 2023a] uses the Phi-2 [Javaheripi et al., 2023] LLM backbone to exceed the performance of larger models.
-LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for better generalizability.
-TinyLLaVA [Zhou et al., 2024] explores the effects of architectural choices and training optimizations, achieving comparable performance to larger models.
-MoE-LLaVA [Lin et al., 2024] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance.
+Google’s Gemini [42] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source.
+Open-source alternatives like LLaVA-1.5 [28] employ advanced compression techniques such as 4/8 bit quantization [11, 38].
+MobileVLM [8] and its improved version, MobileVLM-v2 [9], focus on compact architecture designs and training optimizations for mobile use.
 In most cases, LMM efficiency is enhanced by reducing the size of the backbone of the LMM, but no work has considered the efficiency of the LMM from the perspective of the number of visual tokens.
 2.2 Token Reduction
-The notorious quadratic complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the input sequence length.
-Sparse attention methods such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context.
-Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching.
-One of the main causes of the inefficiency of recent LMMs is their use of a large number of prefix visual tokens that serve as a fixed budget for context [Liu et al., 2023b, Zhu et al., 2023].
-In our study, we present a novel plug-and-play token reduction method based on visual token similarities, which achieves comparable performance with less than one tenth of the original tokens.
+The quadratic complexity of Transformers [44] poses a significant challenge in scaling input sequence length.
+Various approaches try to address this issue.
+Sparse attention methods, e.g., Linformer [46] and ReFormer [22], reduce complexity by limiting attention operations to specific regions rather than the full context.
+Token reduction can also accelerate Transformers [17].
+Methods like [31, 47, 25, 4, 13] focus on reducing the number of tokens within the internal transformer structure, thereby decreasing computational load.
+For instance, token merging [4] employs full attention but progressively reduces tokens in each transformer block by selecting the most representative tokens through bipartite matching.
+However, these uni-modal token reduction methods are not directly applicable to LMMs.
+One of the main inefficiencies in LMMs stems from their use of numerous prefix visual tokens as a fixed context budget [30, 55] (analyzed further in Sec.
+4.2), not from the internal structure of Transformers.
+We discuss the unsuitability of existing uni-modal token reduction methods for LMM acceleration in Sec.
+3.5.
+In our study, we introduce a plug-and-play token reduction method specifically designed for LMMs.
+Our approach, based on visual token similarities, achieves comparable performance while using less than one-fourth of the original tokens.
+The core of our method is a sparsity-based selection mechanism that identifies “anchor” tokens via sparse attention scores within the modality encoder, and is the most crucial design element of PruMerge.
+In parallel to our work, [39] proposes CrossGet, a graph-matching-based algorithm for token matching.
+While both approaches aim to reduce tokens in multimodal contexts, they differ significantly in their methodologies.
+Beyond the token selection module, our token merging module also differs from CrossGet’s graph soft matching.
+Our k-nearest neighbors clustering approach has a time complexity of $O(n)$, which is more computationally efficient compared to CrossGet’s $O(n^{2})$ complexity, thus enhancing scalability and efficiency.
 3 Method: Token Pru-Merging
@@
 3.1 Preliminaries
-Figure 2: The conceptual idea of PruMerge.
-Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3).
-Here $m$ denotes the visual token compression ratio.
-Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024].
+Vision Transformers (ViTs) [12] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [29, 55, 18, 52].
 An input image is divided into a grid of patches and each patch is projected into a token embedding by the ViT.
 In addition to the patch tokens, a class token (i.e., [CLS] token) is computed to aggregate global image information for classification.
-A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [Ba et al., 2016].
-These components work together to improve the model’s capability to understand visual data [Han et al., 2022].
+A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [2].
+These components work together to improve the model’s capability to understand visual data [16].
 In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$.
@@
 The self-attention computes the relevance of each item to other items:
-$\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (3.1)
+$\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (1)
 where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$ and $\mathbf{k}$.
@@
 Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
-$\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (3.2)
+$\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (2)
 The MSA framework allows for simultaneous attention on multiple positions, offering diverse representation subspaces.
@@
 After the self-attention layers is the feed-forward network (FFN), which consists of two linear transformation layers separated by a nonlinear activation function:
-$\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3.3)
+$\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3)
 where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, and $\sigma$ denotes the nonlinear activation function.
@@
 These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which maps the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$.
-The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024].
+The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [52].
 The overall architecture of an LMM is depicted in Figure 1.
-It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022].
+Importantly, the computational cost with these models increases quadratically with the number of input tokens to the LLM [41].
 Mathematically, if there are $N$ tokens in the input, the self-attention mechanism computes a $N\times N$ matrix of attention scores, where each entry in this matrix represents the attention score between a pair of tokens.
@@
 The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
+Figure 2: PruMerge has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3).
+Here $m$ denotes the visual token compression ratio.
 3.2 Adaptive Important Token Selection via Outlier Detection
-(a) Distribution of attention values.
-(b) Image & Tokens Visualizations
-Figure 3: (a) The distribution of attention scores (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens.
-The values in the y-axis are logarithmic values.
-Notably, most of the spatial visual tokens have near-zero attention values with the class token.
-(b) Visualizations of PruMerge and PruMerge+.
-The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022].
+The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [31, 47, 25].
 To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
@@
 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information.
-On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [Radford et al., 2021] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token.
-To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2.
+On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [37] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token.
+To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 2.
 Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3(a).
@@
 Interquartile Range (IQR) Method for outlier detection.
-To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [Boukerche et al., 2020], a statistical technique known for its robustness in outlier detection.
-The essence of the IQR method lies in its capability to establish a boundary or “fence” that delineates the normal range of data.
+To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [5], a statistical technique known for its robustness in outlier detection.
+Its essence lies in its capability to establish a boundary or “fence” that delineates the normal range of data.
 This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1.
@@
 Note that we use the class attention value from the penultimate layer for this calculation.
-As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image.
+As shown in Fig.3(b), the sampled visual tokens show two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image.
 Simpler images such as “Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens.
@@
 Such visualizations align with our visual token sampling design.
-These trends are also observed at the benchmark level; in Table 6, the average token numbers on various benchmarks differ.
+These trends are also observed at the benchmark level; in Tab.6, the average token numbers on various benchmarks differ.
+(a) Distribution of attention values.
+(b) Image & Tokens Visualizations
+Figure 3: (a) Distribution of attention scores (in CLIP-ViT’s penultimate layer) between the [cls] token and visual tokens.
+The y-axis shows logarithmic values.
+Notably, most spatial visual tokens have near-zero attention values with the class token.
+(b) Visualizations of PruMerge and PruMerge+.
 3.3 Token Supplement via Similar Key Clustering
@@
 Here we leverage the self-attention mechanism in ViTs.
-Since the key vector of each patch token already contains information summarized in the self-attention module [Vaswani et al., 2017], the final layer’s key vector serves as the representation.
-And then we use the dot product between keys to calculate which tokens have similar visual information [Bolya et al., 2023]:
-$\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (3.4)
+Since the key vector of each patch token already contains information summarized in the self-attention module [44], the final layer’s key vector serves as the representation.
+And then we use the dot product between keys to calculate which tokens have similar visual information [4]:
+$\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (4)
 which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorized form for the set of all tokens $1,2,\cdots,n$}, where $n$ is the number of input visual tokens.
@@
 The integration of pruned tokens into these clusters is guided by their respective class attentions $\mathbf{a}[i]$, enabling a refined representation of each unpruned token through a weighted sum.
-This procedure is outlined in Algorithm 1.
-3.4 PruMerge+: Bridging the Efficiency-Performance Gap
+This procedure is outlined in Algorithm 1 in the Appendix.
+3.4 PruMerge+
 While PruMerge achieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without drawbacks.
@@
 PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—with minimal performance degradation.
-This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2.
+This improvement is detailed in Algorithm 1 in Appendix, building upon the token selection strategies outlined in Section 3.2.
 A new aspect of PruMerge+ lies in its enhanced token selection process.
@@
 This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing performance losses while still achieving substantial token reduction.
-Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
-1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$.
-The penultimate layer’s output tokens, $\mathbf{Y}=\{\mathbf{y}_{1},\cdots\mathbf{y}_{n}\}$.
-$n$ is the number of input visual tokens.
-2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
-3:Token PruMerge:
-4:Calculate attention between visual token and class token $\mathbf{a}_{[cls]}$ using Equation 3.2.
-5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ indices $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
-6:(Optional for PruMerge+, see Sec. 3.4) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
-7:(Optional for PruMerge+) Spatial-uniformly sample visual tokens with $r_{o}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
-8:(Optional for PruMerge+) Update the selected tokens’ index with $\{i_{1},\cdots,i_{m},i_{m+1},\cdots,i_{2m}\}$.
-9:for $p=\{i_{1},\cdots,i_{m}\}$ do (see Sec. 3.3)
-10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
-11:     Use $\mathbf{y}_{p}$ as cluster center and find the $k$ most similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
-12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
-13:end for
-14:Output a refined stack of visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
-3.5 Discussion
-Distinction from Existing Token Reduction Methods.
-Token reduction methods [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022] have been proposed for accelerating ViT computation speed.
-Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT.
-By contrast, our approach is not specifically designed to increase the efficiency of a ViT.
-It instead aims to improve the overall efficiency of a large multimodal model, where ViT is only one relatively light cost component as shown in Fig.
+3.5 Discussion: Distinction from Existing Uni-Modal Token Reduction Methods
+Uni-Modal token reduction methods [31, 47, 25, 4, 13, 17] primarily focus on accelerating ViT computation speed by progressively reducing token numbers across transformer blocks, thereby lowering the internal ViT computational cost.
+Our approach, however, differs in its primary objective.
+Rather than targeting ViT efficiency, we aim to enhance the overall efficiency of LMMs, where the ViT is just one component with relatively light computational cost, as shown in Fig.
 1.
-There are two benefits of this design.
-First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens.
+This design offers two key advantages: First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens.
 This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details.
-Thus, using previous token merging methods to obtain one refined class token as representation of visual input is not consistent with the literature of large multimodal models.
+Thus, using previous token reduction methods to obtain one refined class token to represent the visual input is not consistent with the literature of LMMs.
 Second, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens.
 This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation.
+Indeed, experiments comparing PruMergewith implementations of uni-modal token reduction methods in LMMs demonstrate significantly better performance for our method (see Sec.
+4.4.4).
 4 Experiments
-We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1.
+We first show the performance of our approach when applied to LLaVA-1.5 in Sec 4.1.
 We then analyze the efficiency improvement by using our PruMerge on LMM in Sec 4.2.
@@
 Table 1: Comparison with large multimodal models on six benchmarks.
-Our proposed approach PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
+Our PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance to the original LLaVA-1.5.
+PT is none standing that only we only use the second phase in LLaVA to train the model.
 Method LLM Res. PT IT VQAv2 SQAI VQAT POPE MME MMB
@@
 LLaVA-1.5 Vicuna-7B 336 558K 665K 78.5 66.8 58.2 85.9 1510.7 64.3
-LLaVA-1.5 + PruMerge Vicuna-7B 336 558K 665K 72.0 68.5 56.0 76.3 1350.3 60.9
-LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.8 68.3 57.1 84.0 1462.4 64.9
+LLaVA-1.5 + PruMerge Vicuna-7B 336 - 665K 72.0 68.5 56.0 76.3 1350.3 60.9
+LLaVA-1.5 + PruMerge+ Vicuna-7B 336 - 665K 76.8 68.3 57.1 84.0 1462.4 64.9
 LLaVA-1.5 Vicuna-13B 336 558K 665K 80.0 71.6 61.3 85.9 1531.3 67.7
-LLaVA-1.5 + PruMerge Vicuna-13B 336 558K 665K 72.8 71.0 58.4 78.5 1428.2 62.3
-LLaVA-1.5 + PruMerge+ Vicuna-13B 336 558K 665K 77.8 71.0 58.6 84.4 1485.5 65.7
+LLaVA-1.5 + PruMerge Vicuna-13B 336 - 665K 72.8 71.0 58.4 78.5 1428.2 62.3
+LLaVA-1.5 + PruMerge+ Vicuna-13B 336 - 665K 77.8 71.0 58.6 84.4 1485.5 65.7
 4.1 Main Results
-We apply our method to LLaVA-1.5 [Liu et al., 2023a], a recent state-of-the-art LMM.
-We further finetune LLaVA-1.5 using LoRA [Hu et al., 2022] for 1 epoch using the LLaVA-1.5 instruction fine-tuning data [Liu et al., 2023a] with our reduced visual tokens.
-We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c].
-As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023].
+We apply our method to LLaVA-1.5 [28], a recent state-of-the-art LMM.
+We further finetune LLaVA-1.5 using LoRA [19] for 1 epoch using the LLaVA-1.5 instruction fine-tuning data [28] with our reduced visual tokens.
+We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [15], ScienceQA [33], TextVQA [40], POPE hallucination bench [24], MME [14], and MMBench [32].
+As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [23] and InstructBLIP [10].
 Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5.
@@
 LLaVA-1.5 w/ PruMerge Vicuna-13B INT4 0.45 29.0 6.8 0.11
-To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [Yuan et al., 2024].
+To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [51].
 Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs.
@@
 It is important to emphasize that the benefits of PruMerge  extend beyond mere efficiency gains.
-Our token reduction strategy can complement other LLM acceleration techniques, such as quantization and factorization [Yuan et al., 2023b].
+Our token reduction strategy can complement other LLM acceleration techniques, such as quantization and factorization [50].
 This orthogonal relationship underscores the versatile potential of PruMerge to contribute to a broader spectrum of efficiency-enhancing strategies.
-4.3 Generalization on Video-LLM
-To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [Lin et al., 2023].
-Video-LLaVA is one of the most popular open-soruced Video-LLMs.
-We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference.
-The outcome of this integration is shown in Table 3.
-Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [Liu et al., 2023a].
-Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively.
-The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance.
-This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs.
-Exploring ways to further capitalize on this redundancy could shape future research directions.
-Table 3: A comparison of different LVMs on video reasoning benchmarks.
-Like Video-LLaVA, ChatGPT-Assistant is employed to evaluate performance.
-The version of ChatGPT is “gpt-3.5-turbo”.
-Note that we directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
+Table 3: Comparison of different LVMs on video reasoning benchmarks.
+Like Video-LLaVA, ChatGPT-Assistant (version ‘gpt-3.5-turbo’) is employed to evaluate performance.
+We directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
 Methods LLM size MSVD-QA MSRVT-QA ActivityNet-QA
@@
 Video-LLaVA + PruMerge+ 7B 71.1 3.9 59.3 3.6 47.7 3.4
+4.3 Generalization on Video-LLM
+To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [26].
+Video-LLaVA is one of the most popular open-soruced Video-LLMs.
+We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference.
+The outcome of this integration is shown in Table 3.
+Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [28].
+Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively.
+The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance.
+This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs.
+Exploring ways to further capitalize on this redundancy could shape future research directions.
 4.4 Ablation Study
@@
 Here we show how our approach performs better than the vanilla visual token sampling strategy, including sequential sampling and spatial sampling.
+Figure 4: Different token sampling strategies.
 LLaVA-PruMerge: Our approach dynamically samples key visual tokens (see Sec.
@@
 The visualization is shown in Figure 4 (b).
-Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apples-to-apples comparison, shown in Figure 4 (c).
-Figure 4: Different token sampling strategies.
-Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h).
+Sequential sampling: We sample $N$ tokens in the flatted visual tokens; e.g., the first 40 tokens are sampled for an apples-to-apples comparison, Fig. 4 (c).
+Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, Fig.
+4 (d-h).
 We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
@@
 Spatial $4\times 4=16$ 66.29
-Table 5: Ablation Studies for Adaptive Important Token Selection (AITS, Sec. 3.2) and Token Supplement (TS, Sec. 3.3). With these modules, the downstream performance can be progressively improved.
+Table 5: Ablation Studies for AITS (Sec. 3.2) and TS (Sec. 3.3). With these modules, the downstream performance can be progressively improved.
 Method LLM SQAI VQAT POPE MME
@@
 LLaVA-PruMerge w. LoRA-FT Vicuna-7B 68.5 56.0 76.3 1350.3
+Table 7: Comparison with SoTA token reduction methods on LMMs. Ratio denotes the proportion of remaining tokens. CrossGet results are directly from its paper.
+Method Ratio VQAv2 SQAI VQAT POPE MME MMB
+LLaVA-1.5 100% 78.5 66.8 58.2 85.9 1510.7 64.3
+LLaVA-1.5 + ToMe 25% 66.0 62.7 56.0 51.0 1385.2 56.9
+LLaVA-1.5 + ATS 25% 66.7 63.0 55.1 57.4 1313.2 54.9
+LLaVA-1.5 + EViT 25% 65.5 64.1 54.2 60.1 1299.3 56.2
+LLaVA-1.5 + PruMerge+ 25% 76.8 68.3 57.1 84.0 1462.4 64.9
+LLaVA-1.5 + CrossGet 50% 77.3 66.7 54.9 83.9 1510.2 64.7
+LLaVA-1.5 + PruMerge+ 50% 77.6 68.5 57.6 85.1 1507.1 64.9
 Note that all the experiments are done via a training-free manner.
 As shown in Table 6, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge.
-Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019].
+Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [40].
 Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details.
-This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(b), where more attentive tokens are distributed on the foreground text in the images.
+This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1 in the Appendix, where more attentive tokens are distributed on the foreground text in the images.
 4.4.2 Effectiveness of Each Module in PruMerge
@@
 4.4.3 Training Analysis: Training-free v.s. Fine-tuning
-Finally, LLaVA-PruMerge can be conducted in either a training-free or fine-tuning manner.
+PruMerge can be conducted in either a training-free or fine-tuning manner.
 With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks.
-As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].
+As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [33], TextVQA [40], POPE [24], and MME [14].
+The results in Table 1 are all taken from the fine-tuned model, LLaVA-PruMerge.
+The results in Table 3 are all taken from the training-free model.
+4.4.4 Comparison to Token Reduction Methods
+To evaluate the effectiveness of our approach, we compare PruMerge+ with SoTA token reduction methods in the context of Large Multimodal Models.
+We utilize the token reduction benchmarking framework from [17] to implement and compare these methods.
+Based on the uni-modal token reduction benchmark, ToMe [4], ATS [13], and EViT [25] are recognized as top performers in uni-modal token reduction.
+We also include CrossGet [39], a concurrent multimodal token reduction method, for comparison.
+Table 7 clearly demonstrates that PruMerge+ significantly outperforms these unimodal token reduction methods on multimodal tasks, supporting our assertion about its superior effectiveness in managing the complexities of multimodal contexts.
+Notably, PruMerge+ outperforms CrossGet [39], a concurrent multimodal token reduction method, under the same reduction ratio.
+Advantages over uni-modal token merging methods: The superior performance of PruMerge+ on multimodal tasks can be attributed to several key factors: (1) Sparsity-based selection: PruMerge leverages the sparsity observed in multimodal encoders (Visual-LLM and Video-LLM), particularly in how attention scores distribute sparsely in the final layer.
+This pattern is less pronounced in unimodal token reduced models where token relevance may not distribute in the same way.
+(2) Efficiency in LMMs: Considering that the bulk of computational demand within LMMs is attributed to the LLM backbone rather than the modality encoder, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens.
+In contrast, unimodal token merging methods focus on the modality encoder’s internal efficiency.
+This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation.
+(3) Accumulation phenomenon: PruMerge capitalizes on the accumulation of sparsity across layers, a characteristic more specific to multimodal models.
+Unimodal token reduction methods that reduce tokens gradually cannot leverage this sparsity effectively.
 5 Conclusion
@@
 We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
-Appendix A Appendix
-A.1 Limitations and Future Work
-Our exploration of LLaVA-PruMerge has two primary limitations.
-First, the visual token compression, while efficient, is not entirely lossless.
-This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge.
-We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps.
-Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings.
-This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2024] framework with a Yi-34B backbone.
-Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
-A.2 Societal Impacts
-In this study, we propose a technique that improves the efficiency of Large Multimodal Models (LMMs), making them more accessible.
-This approach helps to democratize LMMs by lowering deployment costs and hardware barriers, facilitating their use in edge computing.
-However, it does not mitigate the potential misuse of LMMs by malicious actors.
+6 Acknowledgments
+This research is supported by NSF IIS-2525840 (Yan Yan), CNS-2432534 (Yan Yan), ECCS-2514574 (Yan Yan), NIH 1RF1MH133764-01 (Yan Yan) Cisco Research unrestricted gift (Yan Yan), NSF IIS2404180 (Yong Jae Lee), Microsoft Accelerate Foundation Models Research Program (Yong Jae Lee), and Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No.
+2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration, Yong Jae Lee) and (No.
+RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training, Yong Jae Lee).
+This article solely reflects opinions and conclusions of authors and not funding agencies.